OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

Line is given by: \[y - y_1 = m(x-x_1)\] Here m is the slope.. \[y - 5 = 2(x + 1) \implies y - 5 = 2x + 2 \implies 2x - y = -7\]